find the general solution for the equation 2sin^2x-sinx=1. Answer in radians.
Answer by stanbon(75887)

find the general solution for the equation 2sin^2x-sinx=1. Answer in radians.
------------------
2sin^2(x) - sin(x) - 1 = 0
Factor::
(2sin(x)+1)(sin(x)-1) = 0
-------
sin(x) = -1/2 implies x = 6/5(pi) or x = (11/6)pi
-----------------------
sin(x) = 1 implies x = (1/2)pi
-------------
Cheers,
Stan H.
-----------